OpenStudy (anonymous):

Can someone work with me on Simplifying Power of Power: (a^2b^-3/2)^1/3

OpenStudy (callisto):

Using \((a^m)^n = a^{mn}\) \[(a^2b^{-\frac{3}{2}})^\frac{1}{3}\]\[=(a^{2\times \frac{1}{3}}b^{-\frac{3}{2}\times \frac{1}{3}})\]\[=...?\]

OpenStudy (anonymous):

a^2 b^3

OpenStudy (anonymous):

a^2/3 b

OpenStudy (anonymous):

a^2/3 b^1/3